### v3

# CleanAgent: Automating Data Standardization with LLM-based Agents

Abstract.
Data standardization is a crucial part of the data science life cycle. While tools like Pandas offer robust functionalities, their complexity and the manual effort required for customizing code to diverse column types pose significant challenges. Although large language models (LLMs) like ChatGPT have shown promise in automating this process through natural language understanding and code generation, it still demands expert-level programming knowledge and continuous interaction for prompt refinement. To solve these challenges, our key idea is to propose a Python library with declarative, unified APIs for standardizing different column types, simplifying the LLM’s code generation with concise API calls. We first propose Dataprep.Clean, a component of the Dataprep Python Library, significantly reduces the coding complexity by enabling the standardization of specific column types with a single line of code. Then, we introduce the CleanAgent framework integrating Dataprep.Clean and LLM-based agents to automate the data standardization process. With CleanAgent , data scientists only need to provide their requirements once, allowing for a hands-free process. To demonstrate the practical utility of CleanAgent , we developed a user-friendly web application, allowing SIGMOD attendees to interact with it using real-world datasets.
1. Introduction
Data standardization, which is pivotal in the realm of data science, aims to transform heterogeneous data formats within a single column into a unified data format. This crucial data preprocessing step is essential for enabling effective data integration, data analysis, and decision-making.
Example 1. We illustrate the data standardization task in Figure 1. Given the input table , it is obvious that data in the “Admission Date” column and the “Address” column are in different formats, and the data in the cells of the “Admission Date” column includes two different date formats. The goal of data standardization is to unify the data format in each column in , to get the standardized table satisfying the data scientist’s requirements. In Figure 1, the data scientist inputs their requirement to standardize “Admission Date” with the “MM/DD/YYYY HH:MM:SS” format. In the resulting , data in the cells of the “Admission Date” column follows only one date format, i.e., the “MM/DD/YYYY HH:MM:SS” format.
Previously, data scientists heavily relied on libraries such as Pandas (McKinney et al., 2024) for data standardization tasks. Even though Pandas is a powerful tool, achieving data standardization often requires writing hundreds or thousands of lines of code. The standardization process for a single column involves identifying the column type, applying intricate methods such as regular expressions to each cell for validation, and converting each cell into desired formats. Moreover, a table may contain multiple columns, each possibly of a different type, requiring bespoke standardization code for each column type.
Example 2. Still considering the data standardization task in Figure 1. For standardizing “Admission Date” and “Address”, data scientists need to write the datetime standardization code for “Admission Date” and address standardization code for “Address” using regex. An example standardization code for “Address” is shown as follows.
If the input table has other column types such as email and IP addresses, data scientists also need to write standardization code tailored for the new types, which is time-consuming.
Recently, the emerging LLMs have shown the potential to revolutionize this process. By leveraging their natural language understanding and code generation ability, these models could significantly aid data scientists by autonomously generating standardization code in response to conversational prompts. However, this method still necessitates detailed prompt crafting and often involves multi-turn dialogues (Chen et al., 2023) for different column types in the table one by one, which limits the efficiency and practicality of adopting LLMs in the standardization process.
To overcome these limitations, our key idea is to introduce a Python library involving declarative and unified APIs specifically designed for standardizing different column types. This idea lowers the burden of the LLM, as it now only needs to convert natural language (NL) instructions into succinct, declarative API calls instead of lengthy, procedural code. Such an approach simplifies the LLM’s code generation process for data standardization, requiring just a few lines of code.
The pursuit of simplicity, however, introduces two primary challenges. The first challenge (C1) is the design of the declarative and unified APIs for data standardization, ensuring it can effectively reduce the intricacies involved in standardizing specific column types (ideally one line of code per column type). The second challenge (C2) centers on optimizing the interaction between data scientists and LLMs. Our goal is to minimize human involvement, ideally allowing data scientists to input their standardization requirements in one instance, thereby enabling an autonomous and hands-off data standardization process.
To solve C1, we propose the type-specific Clean module in the Dataprep Library, named Dataprep.Clean. By observing and summarizing the common steps of data standardization for specific column types, we design unified APIs clean_type(df, column_name, target_format), where the type represents the desired standardization type, such as date, address, and phone, etc. These unified APIs offer enhanced expressiveness compared to raw Pandas code, reducing the complexity of standardizing specific column types and allowing one to standardize a column with only one line of code.
To solve C2, we propose the CleanAgent framework which automates data standardization with Dataprep.Clean and LLM-based Agents (Xi et al., 2023; Wu et al., 2023). Once users have entered their final goals, the LLM-based Agents can free their hands, autonomously generate reasoning steps, and execute particular tasks. Data scientists only need to input the table being standardized and their requirements, CleanAgent will complete the data standardization process automatically with three steps: annotating the type of each column, generating concise Python code for standardization, and executing the generated Python code.
Example 3. Continuing with Example 1. Given an input table which needs to be standardized and the data scientists’ requirements, the CleanAgent first recognizes that the “Admission Date” column belongs to the date type, and the “Address” column belongs to the address type. According to the column-type annotation results, the CleanAgent generates and executes Python code for standardization by calling the “clean_date” and “clean_address” functions, then returns the standardized table .
We also built a web interface for CleanAgent . It allows the SIGMOD attendees to choose sample data and communicate with CleanAgent for standardization. We provide the demonstration video, which can be found on Youtube.
To summarize, we make the following contributions: (1) We propose Dataprep.Clean, an open-sourced library for reducing the complexity of implementing data format standardization with type-specific standardization functions. (2) We propose CleanAgent , which automates the data standardization process by combining both the advantages of Dataprep.Clean and LLM-based Agents. (3) We deploy CleanAgent as a web application with a user-friendly interface and demonstrate its utility. We also open-sourced the implementation of CleanAgent on Github.
2. Type-Specific Standardization API Design
In this section, we first describe the common steps of data standardization. Then, we introduce the type-specific API design of Dataprep.Clean.
Common Steps of Data Standardization. Inspired by the steps of how human users standardize data cells, we identify three common steps of data standardization. We take the datetime column type as an example to illustrate these steps.
Assume a data scientist is dealing with an datetime column including two records ”Thu Sep 25 10:36:28 2003” and ”1996.07.10 AD at 15:08:56”. The data scientist wants to unify the messy column into a target format ”YYYY-MM-DD hh:mm:ss”.
(1) Split. In the beginning, the data scientist needs to split the datetime string into several single parts, which include one kind of specific information. In our example, the data scientist can get several tokens {’Thu’,’Sep’,’25’,’10’,’36’,’28’,’2003’} from the first record by using space and colon as separators. Each different type has its splitting strategy, which may not always be splitting the string into tokens. For example, the data scientist will split the email string into the username part and the domain part.
(2) Validate. Standardization can only be performed on valid inputs. Thus, the second step should be validation. For example, if the string “little cat” is an instance of the datetime column, this string is invalid, and the data scientist will transform it to a default value like NaN. Intuitively, a valid string indicates that each part of this string after splitting is valid. Usually, the data scientist will recognize and validate each part by their domain knowledge, some corpus or some rules. If every split part is valid, the string is also valid. For instance, the token ’Sep’ can be recognized as a valid representation of a month, and ’2003’ can be recognized as a valid year.
(3) Transform. The last step of standardization is to transform each split part and combine them into the target format. In our example, because the target format is ”YYYY-MM-DD hh:mm:ss”, the month Sep is transformed into number 09 and recombined with other parts to the target "2003-09-25 10:36:28".
The Design of Unified APIs. The goal of our API design is to enable data scientists to complete all the common steps of standardizing one column with a single function call. Simplicity and consistency are considered the principles of API design. The observation of the common steps of data standardization brings the type-specific API design idea. More specifically, we design the API to be in the following form:
clean_type(df, column_name, target_format) |
where clean_type is the function name, type represents the type of the current column. The first argument df represents the input DataFrame, the second argument column_name is the column being standardized, and the third argument target_format is the target standardization format users specified. Our API design is flexible and extensible, which makes it convenient for users to add their standardization functions for new data types. Currently, we have 142 standardization functions in Dataprep.Clean, each handles one data type.
3. CleanAgent Workflow
In this section, we first introduce the basic structure of LLM-based agents. Then, we describe the CleanAgent workflow constructed by four agents. The automatic data standardization process can be completed by the cooperation of the four agents in CleanAgent .
Basic Structure of LLM-based Agents. According to the previous surveys on LLM-based Agent (Xi et al., 2023), we conclude the basic structure of LLM-based Agents shown in Figure 2. An LLM-based agent includes four main components: (1) a backbone LLM used to generate replies for input prompts, (2) a memory used to store historical conversation messages, (3) a system message defining the role of the agent, and (4) a set of external tools which can be called by the LLM-based agent to complete specific tasks, such as web searching, code execution, etc.
Detailed Workflow. The detailed workflow of CleanAgent is shown in Figure 3. The CleanAgent is composed of four agents, including a Chat Manager, a Column-type Annotator, a Python Programmer, and a Code Executor. They can communicate with each other and automatically complete the data standardization process by cooperation. As mentioned in Figure 2, each agent has its own memory to store the historical conversational messages between it and other agents. Note that the memory of the Chat Manager is uniquely comprehensive, encompassing the entire historical conversational messages from all agents within the CleanAgent system. This extensive memory enables every agent in the CleanAgent to generate responses that are informed by the complete historical messages.
The input of CleanAgent includes a table that needs to be standardized. Data scientists can also input extra requirements such as “the format of the date type column should be MM/DD/YYYY”. By receiving the input table and data scientists’ extra requirements, CleanAgent stores this information into the Chat Manager’s memory and then completes the data standardization process. The Chat Manager delivers messages in its memory to the Column-type Annotator(① in Figure 3). Then, The Column-type Annotator receives the table information and leverages an LLM to annotate the type of each column in the input table. If the The Column-type Annotator cannot figure out the specific type of one column, the Column-type Annotator outputs “I do not know”. The annotation result is returned to the Chat Manager and stored in the Chat Manager’s memory (② in Figure 3). Thirdly, the Python Programmer receives historical messages from the Chat Manager including the column-type annotation results (③ in Figure 3), picks up the corresponding clean functions, and generates Python code for the data standardization process. The generated Python code is also returned to the Chat Manager and stored in the Chat Manager’s memory (④ in Figure 3). Finally, the Code Executor receives historical messages from the Chat Manager including the column-type annotation results and the generated Python code (⑤ in Figure 3), then executes the generated Python code. If the generated code executes without errors, the standardized table is returned; otherwise, the error message is returned to the Chat Manager and stored in its memory (⑥ in Figure 3). Then, CleanAgent will retry the whole workflow until it can complete the data standardization process successfully.
4. Demonstration Scenarios
The demonstration setup requires a laptop that can connect to the Internet so the audience can use CleanAgent smoothly with the LLM API services (we used OpenAI’s GPT). If the conference Internet fails, a mobile hotspot (established via cell phone) can also be used for running CleanAgent .
Figure 4 shows the user interface of CleanAgent . As area ① shows, users must first upload a CSV file that needs to be cleaned. Then CleanAgent shows the basic information of the uploaded file (number of rows and number of columns). If the users can click the “Start Standardization” button to start the data standardization process by want CleanAgent .
After clicking the “Start Standardization” button, as area ② shows, the User_Proxy generates three detailed steps to complete the data standardization task. Firstly, the Column-type Annotator receives messages from the Chat Manager, annotates and outputs the type of each column, as area ③ shows. Then, the Python Programmer picks up standardization functions from Dataprep.Clean based on the type of each column, and write proper Python code using the standardization functions, as area ④ shows. Thirdly, the Code Executor executes the Python code by the Python Programmer and collects the execution messages, as area ⑤ shows. If the Code Executor gets an error message when executing generated Python code, the error message is sent to the Chat Manager and becomes part of the prompt of the next try. If the Code Executor gets the message of successful execution, CleanAgent will report that the data standardization is completed, as area ⑥ shows. Moreover, users can click the “Show Cleaned Table” button to check whether the standardized table matches their requirements. If so, users can download the standardized table directly. Otherwise, users can input their extra requirements with natural language, and CleanAgent will start a new data standardization process accordingly.
5. Conclusion
In this demo paper, we proposed CleanAgent to automate the data standardization process with Dataprep.Clean and LLM-based Agents. We implemented CleanAgent as a web application to visualize the conversations among agents. Other tasks in the data science life cycle, such as data cleaning and data visualization, can also be completed by LLM-based agents (Xue et al., 2023). In the future, it is promising that the data science life cycle can be automatically planned and completed by LLM-based agents’ cooperation.